# T-MAC: CPU Renaissance via Table Lookup for Low-Bit LLM Deployment on Edge

### v1

@@ Abstract.
Our LUT-based kernels scale linearly to the weight bit-width. Evaluated on low-bit Llama and BitNet models, T-MAC demonstrates up to 4$\times$ increase in throughput and 70% reduction in energy consumption compared to llama.cpp. For BitNet-b1.58-3B, T-MAC delivers a token generation throughput of 30 tokens/s with a single core and 71 tokens/s with eight cores on M2-Ultra, and 11 tokens/s on lower-end devices like Raspberry Pi 5, which significantly exceeds the adult average reading speed. T-MAC with LUT-based computing paradigm, paves the way for the practical deployment of low-bit LLMs on resource-constrained edge devices without compromising computational efficiency. The system is open-sourced at https://github.com/microsoft/T-MAC.
submissionid: 1060
* Work done during an internship at Microsoft Research
@@ 1. Introduction
We implemented T-MAC on the pervasively available CPU processors of edge devices, even a Raspberry Pi. We find T-MAC makes the LLM inference speed on a CPU comparable or even higher to the GPU on the same device, mainly because T-MAC has no conversion cost and the total operation reduction by table lookup. This paper is thus the first to provide a practical solution for deploying LLMs on edge devices using the widely available CPUs, without relying on GPUs. We plan to open-source T-MAC. We evaluated T-MAC performance on typical edge devices, including Apple M2 Ultra, Jetson AGX Orin, Surface Book 3, and Raspberry Pi 5. The T-MAC kernel speedup can reach up to 6.6$\times$ and an average of 3.6$\times$ compared to the SOTA on the CPU by llama.cpp (llama.cpp, ). The e2e LLM inference speedup achieves 2.8$\times$ speedup for Llama-2-7B-2bit model (gguf-models, ). The inference performance can reach 11.1 tokens/s even on a Raspberry Pi for BitNet-b1.58-3B model (wang2023bitnet, ). T-MAC is also energy efficient, reducing 60-70% energy compared to llama.cpp.
@@ 3.1. T-MAC Algorithm
Figure 2 and and Alg. 1 shows the T-MAC design. During the offline preparation stage (line 29 to 35), a $n$-bit weight matrix is decomposed into $n$ one-bit matrices. Since one bit can only represent two values, for a group with $g$ bits, the possible permutations are only $2^{g}$. The permutations can be precomputed with each $g$-width vector of the activation and saved in a table. A $g$-bit group in the weight is thus an index to look up the table for the precomputed results. Therefore, a table in T-MAC is defined to save the results of a $[1,g]\times[g,2^{g}]$ sub-matrix multiplication, and the table size is $[1,2^{g}]$. During the offline stage, a tile in the one-bit matrix will be saved continuously in memory to facilitate fast loading. Same as the tiling of normal matrix multiplication, a tiling here is also to improve data locality and cache utilization during LUT.
@@ 4. Implementation
Instruction Set Look-up Fast Aggregation
NEON vqtbl1q_u8 vrhaddq_u8
AVX2 _mm256_shuffle_epi8 _mm256_avg_epu8
Table 1. Hardware Intrinsics for Look-up and Aggregation
@@ 5.1. Evaluation Setup
To evaluate the performance of T-MAC, we conduct extensive benchmarks using real-word low-bit LLMs and scenarios. For the kernel performance benchmark, we select matrix shapes derived from the Llama-2-7B and Llama-2-13B models, ensuring our evaluation reflects the practical demands. To conduct an end-to-end throughput test, we employed actual quantized models to demonstrate the practical efficacy of T-MAC across different bit-width configurations. Specifically, we employ 4-bit,3-bit,2-bit and 1-bit quantized Llama models, and also 1-bit and 1.58bit BitNet models that are trained from scratch. The 4-bit Llama models are from GPTQ (frantar2022gptq, ). The 3-bit and 2-bit Llama models are from BitDistiller (du2024bitdistiller, ). The 1-bit Llama models are from OneBit (xu2024onebit, ).
@@
We compared the performance of T-MAC with llama.cpp, a state-of-the-art implementation for LLM deployment on edge devices. We chose llama.cpp as the baseline for several compelling reasons. Firstly, llama.cpp represents the cutting-edge in LLM deployment on edge devices, featuring highly optimized kernel implementations tailored to each hardware platform. Its versatility and robust performance make it an ideal benchmark for assessing the efficacy of new methodologies. Additionally, llama.cpp is implemented in plain C/C++ without any dependencies, ensuring maximum compatibility and efficiency across diverse hardware configurations. For kernel performance benchmarks, we utilized the optimized kernels provided by llama.cpp as the baselines on the respective hardware devices. In our end-to-end throughput evaluations, we integrate the LUT-based kernels from T-MAC to llama.cpp and compare it with original llama.cpp.
@@ 5.6. Error Analysis
5.7. Compared with GPU
GPUs are widely used in LLM deployments. We compare T-MAC on CPU with llama.cpp on GPU to illustrate the efficiency of T-MAC. llama.cpp is the state-of-the-art LLM inference framework on both CPU and GPU.

### v2

@@ Abstract.
Our LUT-based kernels scale linearly to the weight bit-width. Evaluated on low-bit Llama and BitNet models, T-MAC demonstrates up to 4$\times$ increase in throughput and 70% reduction in energy consumption compared to llama.cpp. For BitNet-b1.58-3B, T-MAC delivers a token generation throughput of 30 tokens/s with a single core and 71 tokens/s with eight cores on M2-Ultra, and 11 tokens/s on Raspberry Pi 5. T-MAC with LUT-based computing paradigm, paves the way for the practical deployment of low-bit LLMs on resource-constrained edge devices without compromising computational efficiency. The system is open-sourced at https://github.com/microsoft/T-MAC.
Large Language Models, General Matrix Multiplication, Low-bit Quantization, Computing Algorithms
submissionid: 1060journalyear: 2025copyright: acmlicensedconference: Twentieth European Conference on Computer Systems; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Twentieth European Conference on Computer Systems (EuroSys ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3689031.3696099isbn: 979-8-4007-1196-1/25/03ccs: Computing methodologies Vector / streaming algorithmsccs: Computing methodologies Parallel algorithmsccs: Computing methodologies Natural language generation
@@ 1. Introduction
We implemented T-MAC on the pervasively available CPU processors of edge devices, even a Raspberry Pi. We find T-MAC makes the LLM inference speed on a CPU comparable or even higher to the GPU on the same device, mainly because T-MAC has no conversion cost and the total operation reduction by table lookup. This paper is thus the first to provide a practical solution for deploying LLMs on edge devices using the widely available CPUs, without relying on GPUs. We evaluated T-MAC performance on typical edge devices, including Apple M2 Ultra, Jetson AGX Orin, Surface Book 3, and Raspberry Pi 5. The T-MAC kernel speedup can reach up to 6.6$\times$ and an average of 3.6$\times$ compared to the SOTA on the CPU by llama.cpp (llama.cpp, ). The e2e LLM inference speedup achieves 2.8$\times$ speedup for Llama-2-7B-2bit model (gguf-models, ). The inference performance can reach 11.1 tokens/s even on a Raspberry Pi for BitNet-b1.58-3B model (wang2023bitnet, ). T-MAC is also energy efficient, reducing 60-70% energy compared to llama.cpp.
@@ 3.1. T-MAC Algorithm
Figure 2 and and Alg. 1 shows the T-MAC design. During the offline preparation stage (line 29 to 35), a $n$-bit weight matrix is decomposed into $n$ one-bit matrices. Since one bit can only represent two values, for a group with $g$ bits, the possible permutations are only $2^{g}$. During the online stage, the permutations can be precomputed with each group of activation of shape $[1,g]$ and saved in a table. A $g$-bit group in the weight is thus an index to look up the table for the precomputed results. Therefore, a table in T-MAC is defined to save the results of a $[1,g]\times[g,2^{g}]$ sub-matrix multiplication, and the table size is $[1,2^{g}]$. During the offline stage, a tile in the one-bit matrix will be saved continuously in memory to facilitate fast loading. Same as the tiling of normal matrix multiplication, a tiling here is also to improve data locality and cache utilization during LUT.
@@
To illustrate with an example, given g=4, for an activation (A1,A2,A3,A4) of shape $[1,4]$ and 1-bit weights of shape $[4,M]$, the activation will be precomputed online into a LUT of shape $[1,16]$, containing elements from -A1-A2-A3-A4 to A1+A2+A3+A4. By grouping every 4 weights together, the weight vector of $0$ will lookup -A1-A2-A3-A4 and 0101 will lookup -A1+A2-A3+A4.
@@ 4. Implementation
Instruction Set Look-up Fast Aggregation
NEON vqtbl1q_u8 vrhaddq_u8
AVX2 _mm256_shuffle_epi8 _mm256_avg_epu8
Table 1. Hardware Intrinsics for Look-up and Aggregation
@@
Determine the size of on-chip LUT
The number of on-chip LUTs is tuned for each hardware, to make sure on-chip memory can be fully utilized and LUTs won’t be swapped out for the tile. Larger on-chip memory allows more LUTs to reside, enabling more intermediate results to be aggregated before writing back. Register spill caused by excessive on-chip LUTs can lead to overhead.
$g$ is also determined by the on-chip memory size and instruction throughput. LUT for $g=4$ exactly fits into one register for ARM.TBL/AVX2.PSHUF. A larger $g$, such as $g=5$, requires two registers and slower ARM.TBL2/AVX512.PSHUF.
@@ 5.1. Evaluation Setup
To evaluate the performance of T-MAC, we conduct extensive benchmarks using real-word low-bit LLMs and scenarios. For the kernel performance benchmark, we select matrix shapes derived from the Llama-2-7B and Llama-2-13B models, ensuring our evaluation reflects the practical demands. To conduct an end-to-end throughput test, we employed actual quantized models to demonstrate the practical efficacy of T-MAC across different bit-width configurations. Specifically, we employ 4-bit,3-bit,2-bit and 1-bit quantized Llama models, and also 1-bit and 1.58bit BitNet models that are trained from scratch. Ternary weights in 1.58bit BitNet are interpreted as 2-bit and decomposed into two 1-bit matrices. The 4-bit Llama models are from GPTQ (frantar2022gptq, ). The 3-bit and 2-bit Llama models are from BitDistiller (du2024bitdistiller, ). The 1-bit Llama models are from OneBit (xu2024onebit, ).
@@
We compared the performance of T-MAC with llama.cpp (version b2794, realesed on May 2024), a state-of-the-art implementation for LLM deployment on edge devices. We chose llama.cpp as the baseline for several compelling reasons. Firstly, llama.cpp represents the cutting-edge in LLM deployment on edge devices, featuring highly optimized kernel implementations tailored to each hardware platform. Its versatility and robust performance make it an ideal benchmark for assessing the efficacy of new methodologies. Additionally, llama.cpp is implemented in plain C/C++ without any dependencies, ensuring maximum compatibility and efficiency across diverse hardware configurations. For kernel performance benchmarks, we utilized the optimized kernels provided by llama.cpp as the baselines on the respective hardware devices. In our end-to-end throughput evaluations, we integrate the LUT-based kernels from T-MAC to llama.cpp and compare it with original llama.cpp.
We also compared the performance of T-MAC with llama.cpp (BLAS). llama.cpp uses Accelerate on M2-Ultra and OpenBLAS on the other platforms. llama.cpp (BLAS) is slower for mpGEMV but faster for mpGEMM compared to llama.cpp’s highly optimized mixed-precision implementation. Therefore, T-MAC is measured against BLAS only for mpGEMM.
@@
T-MAC demonstrates a significant advantage at 3-bit precision. This can be attributed to the inefficiencies in current techniques for handling 3-bit weights. Weights decoding is typically executed using SHIFT and AND instructions, which require weights being aligned with the width of bytes. Since 8 is indivisible by 3, this decoding process is notably inefficient. llama.cpp attempts to optimize it by separately packing 2 bits and the remaining 1 bit, but it still results in significant overhead. In contrast, T-MAC avoids this problem by individually computing the results of each bit.
@@ 5.6. Error Analysis
5.7. Compared with GPU/NPU
GPUs are widely used in LLM deployments. We compare T-MAC on CPU with llama.cpp on GPU to illustrate the efficiency of T-MAC. llama.cpp is the state-of-the-art LLM inference framework for both CPU and GPU on edge devices.
@@ 5.7. Compared with GPU/NPU
Device Total CPU Cores Used Cores GPU NPU
Surface 12 Oryon, 3.8 GHz 4 Adreno X1-85 Hexagon
Laptop 7 4.6 TFLOPS 45 TOPS
OnePlus 12 1 Cortex-X4, 3.2 GHz 1 Adreno 750 4.6 TFLOPS Hexagon 15 TOPS
5 Cortex-A720, 3.0 GHz 3
2 Cortex-A520, 2.3 GHz 0
Jetson 8 Cortex A78AE, 2.0 GHz 6 Ampere GA10B -
Orin NX 50 TOPS
Table 6. Detailed CPU/GPU/NPU specifications of the tested platforms. ”Used Cores” refers to the number of CPU cores we used to fully utilize the memory bandwidth.
Device Model T-MAC (CPU) llama.cpp (CPU) llama.cpp (GPU) NPU
Surface Laptop 7 Llama-2-7B-4bit 21.63 10.64 - 10.40
Llama-2-7B-2bit 31.83 9.39 10.40*
OnePlus 12 Llama-2-7B-4bit 10.19 8.24 1.60 11.30
Llama-2-7B-2bit 16.62 6.95 1.72 11.30*
Jetson Orin NX Llama-2-7B-4bit 7.53 3.97 14.76 -
Llama-2-7B-2bit 11.41 3.20 7.94
Table 7. Token generation tokens/s of T-MAC vs GPU/NPU for Llama-2-7B-4bit/2bit on three devices. The 2-bit performance of NPUs is deduced from 4-bit and marked with ”*”.
@@
In addition to its power efficiency, T-MAC also demonstrates superior performance over GPU/NPUs across widely used platforms. We further evaluate T-MAC on three devices: Surface Laptop 7, OnePlus 12 and Jetson Orin NX. The full specifications of equipped CPU/GPU/NPUs are detailed in Table 6. We utilize the minimum number of CPU cores that can fully leverage the memory bandwidth and nearly achieve optimal performance. GPU evaluations use the llama.cpp CUDA backend for the NVIDIA GPU and OpenCL backend for the Qualcomm GPU. The performance of NPUs are sourced from official data released by Qualcomm via Qualcomm AI Hub (qualcomm-ai-hub, ).
With T-MAC, CPUs can achieve much higher computation throughput and fully exploit the memory bandwidth. GPU/NPUs, which are also memory-bound for GEMV during token generation, share unified memory with CPUs on most edge devices. As demonstrated in Table 7, T-MAC achieves significant speedup for Llama-2-7B-2bit on all three devices. Specifically, T-MAC achieves 3$\times$ speedup on Surface Laptop 7 over the NPU with only 4 out of the total 12 CPU cores, 1.5$\times$ speedup over the NPU on OnePlus 12, and 1.4$\times$ speedup over the Ampere GPU on the Jetson Orin NX. Even for Llama-2-7B-4bit, T-MAC maintains 2.1$\times$ speedup on Surface Laptop 7. Notably, on OnePlus 12, T-MAC demonstrates a substantial speedup of 6.4$\times$ and 9.7$\times$ over the Adreno GPU for 4-bit and 2-bit respectively.
In summary, T-MAC, leveraging widely available CPUs, delivers a notable performance advantage over GPUs, and even NPUs specially designed for AI workloads. This makes T-MAC a practical solution for LLM deployment on edge devices.